OpenStudy (anonymous):

A rectangular football pitch has its length equal to twice its width and a perimeter of 360m.Find its length and width.

terenzreignz (terenzreignz):

What's the formula for the perimeter of a rectangle?

OpenStudy (anonymous):

P=(2)(l+w)

terenzreignz (terenzreignz):

That's good :) We also know that the perimeter is 360, so you can go ahead and replace P with 360... \[\large 360 = 2(l + w)\]

terenzreignz (terenzreignz):

Now, the length... is equal to twice the width. So what does that tell you about the relationship between l and w?

OpenStudy (anonymous):

yes.. but i dont know what is the value for w :D

terenzreignz (terenzreignz):

Well, you know 'length' is TWICE the width. So you just have to translate that into math-language :P l = 2w

terenzreignz (terenzreignz):

Does that make sense? Twice the width.. that means multiplying the width by 2.

OpenStudy (anonymous):

ok.. wait

OpenStudy (anonymous):

have to look for the value of width

OpenStudy (anonymous):

answer?

OpenStudy (souvik):

put l=2w in this equation \[360=2(l+w)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the answer is 60

OpenStudy (souvik):

good ...now calculate l

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks

OpenStudy (souvik):

my pleasure...:)

OpenStudy (anonymous):

i need help for this

OpenStudy (anonymous):

hello

OpenStudy (souvik):

which one?no11?

OpenStudy (anonymous):

11 and all numbers in test 2 please... i'm so tired na

OpenStudy (anonymous):

need help so bad

OpenStudy (souvik):

11.a) multiply first equation with 4 and multiply second one with 3 and add them..